four times the smallest of three consecutive add integers is five more than three times the largest. What is the sum of these consecutive odd integers?
Answer by Mathtut(3670)

let integers be a ,a+2 ,and a+4
:
4a=3(a+4)+5
:
4a=3a+12+5
:
a=17
a+2=19
a+4=21
:
sum = 17+19+21=57